# Mixed Berndt-Type Integrals and Generalized Barnes Multiple Zeta Functions

### v2

@@ 4 Mixed Berndt-Type Integrals via Hyperbolic Series
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+$ $\left(-1\right)^{\mu}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=-\left(\displaystyle\frac{1-i}{2}\pi\right)^{2\mu+1}\bar{S}_{2\mu,2}\left(\displaystyle\frac{\pi}{2}\right),$ (4.39)
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x+i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+\left(-1\right)^{\mu}i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=4\left(1-i\right)^{2\mu-1}\pi^{2\mu}\left[\mu S_{2\mu-1,1}\left(2\pi\right)-\pi\mathrm{DS}_{2\mu,2}\left(2\pi\right)\right].$ (4.41)
@@ 5 Evaluations of Hyperbolic Summations via Jacobi Functions
Now, we evaluate the the reciprocal hyperbolic series of Ramanujan type
@@
by applying the Fourier series expansions and the Maclaurin series expansions of a few Jacobi elliptic functions. Let $x,y$ and $z$ satisfy the relations in (2.14).
@@
Now, we need to compute the series $S_{2p-1,1}^{{}^{\prime}}\left(y\right)$. According to Lemma 3.1 or Lemma LABEL:sn_coefficient_symmetry, there are polynomials where ${\rm q}_{2n-1}(x)\in\mathbb{Z}[x]$ such that the Maclaurin series of ${\rm sn}(u)$ has the form
@@
From [8], we record the Fourier series expansions of Jacobi elliptic functions ${\rm sn}(u)$
@@
Applying $q\equiv q\left(x\right)=e^{-y}$, then
${\rm sn}(u)$ $=\displaystyle\frac{2\pi}{\mathrm{K}k}\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{q^{n+\tfrac{1}{2}}\left(2n+1\right)^{2j+1}}{1-q^{2n+1}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j+1}}{\left(2j+1\right)!}}$
$=\displaystyle\frac{\pi}{\mathrm{K}k}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2\mathrm{K}}\right)^{2j+1}}{\left(2j+1\right)!}}S_{2j+1,1}^{\prime}\left(y\right).$ (5.46)
Comparing the coefficients of $u^{2m+1}$ in (5.44) and (5), we deduce
@@
Differentiating (5.47) yields
@@
Using the relation $dx/dy=-x(1-x)z^{2}$ (see [5, P. 120, Entry. 9(i)]), we obtain
@@
Finally, combining (5.43), (5.47), (5.49) and (5) yields the desired result with an elementary calculation. ∎
By Mathematica we can compute the following easily using the formulas in Theorem 5.1.
@@
where $E$ is the complete elliptic integrals of the second kinds, see Section 2. Applying (5.44), we deduce the power series expansion of the function $\mathrm{sn}^{2}(u)$
@@
Differentiating (5.53) gives
@@
From [23, Cor.1], we also obtain
@@ 6 Evaluations of Mixed Berndt-type Integrals
Assuming that $x_{0},y_{0}$ and $z_{0}$ are all known. For $\Omega\left(x_{0},\exp\left(-y_{0}\right),z_{0},z_{0}^{\prime}\right)=0$, considering the transformation (3.18)
@@
According to Lemma LABEL:sn_coefficient_symmetry, we know that $\text{q}_{2p-1}(x)=\displaystyle\sum^{p-1}_{j=0}a_{j}x^{j}$. Applying Theorem 5.1 and transformation (6.58), one obtains
@@
If letting $x_{0}=1/2$ in (6.61), then $y_{0}=\pi$ by (2.14). And the coefficients before $z_{0}^{2p+2},z_{0}^{2p+1}z_{0}^{\prime}$ are all rational numbers. when $x_{0}=1/2$, the coefficient of $z_{0}^{2p+1}z_{0}^{\prime}$ is
@@
In fact, Lemma LABEL:sn_coefficient_symmetry indicates that $a_{j}=a_{p-1-j}$, for $j=0,\ldots,p-1$, Thus,
@@
$=a_{j}\left\{\left(\sqrt{2}-1\right)^{4j+2p-2}+\left(\sqrt{2}+1\right)^{2p-2}\left(\sqrt{2}-1\right)^{4\left(p-1-j\right)}\right\}$
$=a_{j}\left\{\left(\sqrt{2}-1\right)^{4j+2p-2}+\left(\sqrt{2}+1\right)^{4j+2p-2}\right\}$
$=a_{j}\displaystyle\sum_{l=0}^{4j+2p-2}{\left(\begin{array}[]{c}4j+2p-2\\ l\\ \end{array}\right)}\left[\left(-1\right)^{l}+1\right]2^{2j+p-1-\tfrac{l}{2}}$ (6.65)
$=2a_{j}\displaystyle\sum_{m=0}^{2j+p-1}{\left(\begin{array}[]{c}4j+2p-2\\ 2m\\ \end{array}\right)}2^{2j+p-1-m}\in\mathbb{Q}\quad(j=1,\ldots,[(p-1)/2]).$ (6.68)
Further, for $j=0,\ldots,p-1$, we obtain
@@
$\quad\quad=\displaystyle\frac{\left(-1\right)^{p}}{2^{3p+1}}\sqrt{2}\left(1+\sqrt{2}\right)^{2p-4}\displaystyle\sum_{j=0}^{p-1}{a_{j}}\left\{\left(2+\sqrt{2}\right)p-\left(4j+2\right)\right\}\left(\sqrt{2}-1\right)^{4j-2}.$ (6.76)
According to $a_{j}=a_{p-1-j},j=0,\ldots,p-1$, we have
$\sqrt{2}\left(1+\sqrt{2}\right)^{2p-4}$$\left\{\begin{array}[]{c}a_{j}\left\{\left(2+\sqrt{2}\right)p-2\left(2j+1\right)\right\}\left(\sqrt{2}-1\right)^{4j-2}+\\ a_{p-1-j}\left\{\left(2+\sqrt{2}\right)p-2\left(2p-2j-1\right)\right\}\left(\sqrt{2}-1\right)^{4p-4j-6}\\ \end{array}\right\}$ (6.79)
$=a_{j}\left\{\begin{array}[]{c}\left\{\left(2\sqrt{2}+4\right)p-2\sqrt{2}\left(2j+1\right)\right\}\left(\sqrt{2}+1\right)^{2p-4j-2}+\\ \left\{\left(2\sqrt{2}+4\right)p-2\sqrt{2}\left(2p-2j-1\right)\right\}\left(\sqrt{2}-1\right)^{2p-4j-2}\\ \end{array}\right\}$ (6.82)
$=a_{j}\left\{\begin{array}[]{c}2\sqrt{2}\left(2j+1\right)\left\{\left(\sqrt{2}-1\right)^{2p-4j-2}-\left(\sqrt{2}+1\right)^{2p-4j-2}\right\}\\ 2\sqrt{2}p\left\{\left(\sqrt{2}-1\right)^{2p-4j-1}+\left(\sqrt{2}+1\right)^{2p-4j-1}\right\}\\ \end{array}\right\}$ (6.85)
$=4a_{j}\left\{\begin{array}[]{c}p\displaystyle\sum_{l=0}^{p-2j-1}{\left(\begin{array}[]{c}2p-4j-1\\ 2l\\ \end{array}\right)2^{p-2j-l}}\\ -\left(2j+1\right)\displaystyle\sum_{l=0}^{p-2j-2}{\left(\begin{array}[]{c}2p-4j-2\\ 2l+1\\ \end{array}\right)2^{p-2j-1-l}}\\ \end{array}\right\}\in\mathbb{Q},$ (6.92)
where $j=1,\ldots,[(p-1)/2]$.
Combining (6.69) and (6.92) gives
@@
When $x_{0}=1/2$ then $y_{0}=\pi,\ z_{0}=\displaystyle\frac{\Gamma^{2}}{2\pi^{3/2}}$ and $z_{0}^{\prime}=\displaystyle\frac{4\pi^{1/2}}{\Gamma^{2}}$ by (2.14), then substitute it into (6.93) yields
@@
In Theorem 4.1, letting $a=4m-2$ and $4m$, we have
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{8}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{33\Gamma^{20}}{2097152\pi^{6}}-\displaystyle\frac{63\Gamma^{16}}{262144\pi^{4}}.$
@@
×$\quad\quad\times\left[\begin{array}[]{c}z_{0}\left(1+\sqrt{x_{0}}\right)\displaystyle\lim_{x=4\sqrt{x_{0}}/(1+x_{0}+2\sqrt{x_{0}})}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}\\ +2p\displaystyle\frac{1+x_{0}-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\mathrm{Q}_{2p-2}\left(\displaystyle\frac{1+x_{0}-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\right)\left\{\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}\right\}\\ \end{array}\right]$ (6.105)
@@
×$\quad\quad\times\left\{\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}\right\}+\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{2}{z_{0}}^{2p+2}}{2^{2p-2}}\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-4}$
×$\quad\quad\times\displaystyle\lim_{x=4\sqrt{x_{0}}/(1+x_{0}+2\sqrt{x_{0}})}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}.$ (6.106)
@@
$\mathrm{Q}_{2p-2}\left(x\right)=\displaystyle\sum_{j=0}^{p-2}{b_{j}x^{j}}\quad(b_{j}\in\mathbb{Q}),$ (6.108)
@@
$=b_{j}\left\{\left(\sqrt{2}-1\right)^{4j-2p+4}+\left(\sqrt{2}+1\right)^{4j-2p+4}\right\}$
$=b_{j}\displaystyle\sum_{l=0}^{4j-2p+4}{\left(\begin{array}[]{c}4j-2p+4\\ l\\ \end{array}\right)}\left[\left(-1\right)^{l}+1\right]2^{2j-p+2-\tfrac{l}{2}}$ (6.111)
$=2b_{j}\displaystyle\sum_{m=0}^{2j-p+2}{\left(\begin{array}[]{c}4j-2p+4\\ 2m\\ \end{array}\right)}2^{2j-p+2-m}\in\mathbb{Q}\quad(j=0,\ldots,p-2).$ (6.114)
@@
$=\displaystyle\frac{p}{4}\left[\left(\sqrt{2}+1\right)^{2p-4j-5}-\left(\sqrt{2}-1\right)^{2p-4j-5}\right]+\displaystyle\frac{\sqrt{2}}{2}\left(j+1\right)\left[\left(\sqrt{2}-1\right)^{2p-4j-4}-\left(\sqrt{2}+1\right)^{2p-4j-4}\right],$
$=\displaystyle\frac{p}{4}\displaystyle\sum_{l=0}^{2p-4j-5}{\left(\begin{array}[]{c}2p-4j-5\\ l\\ \end{array}\right)\left[1-\left(-1\right)^{l}\right]2^{p-2j-\tfrac{l+5}{2}}}+\displaystyle\frac{j+1}{2}\displaystyle\sum_{l=0}^{2p-4j-4}{\left(\begin{array}[]{c}2p-4j-4\\ l\\ \end{array}\right)\left[\left(-1\right)^{l}-1\right]2^{p-2j-\tfrac{l+3}{2}}}$ (6.127)
$=\displaystyle\frac{p}{2}\displaystyle\sum_{l=0}^{p-2j-3}{\left(\begin{array}[]{c}2p-4j-5\\ 2l+1\\ \end{array}\right)2^{p-2j-3-l}}-\left(j+1\right)\displaystyle\sum_{l=0}^{p-2j-2}{\left(\begin{array}[]{c}2p-4j-4\\ 2l+1\\ \end{array}\right)2^{p-2j-2-l}}\in\mathbb{Q},$ (6.132)
where $j=0,\ldots,[(p-2)/2]$.
Combining (6.115) with (6.122), we derive that
@@
In Theorem 4.2, letting $a=4m$ and $24m-$ 2, we have

### v4

@@ 3 Mixed Berndt-Type Integrals via Hyperbolic Series
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu+2}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+$ $\left(-1\right)^{\mu+1}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu+2}\left(\sin x-i\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=-\left(\displaystyle\frac{1-i}{2}\pi\right)^{2\mu+3}\bar{S}_{2\mu+2,2}\left(\displaystyle\frac{\pi}{2}\right),$ (3.41)
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu+2}\left(\sinh x+i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}+\left(-1\right)^{\mu+1}i\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{2\mu+2}\left(\sinh x-i\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=4\left(1-i\right)^{2\mu+1}\pi^{2\mu+2}\left[(\mu+1)S_{2\mu+1,1}\left(2\pi\right)-\pi\mathrm{DS}_{2\mu+2,2}\left(2\pi\right)\right].$ (3.43)
@@ 4 Evaluations of Hyperbolic Summations via Jacobi Functions
Now, we evaluate the reciprocal hyperbolic series of Ramanujan type
@@
by applying the Fourier series expansions and the Maclaurin series expansions of certain Jacobi elliptic functions. Let $x,y$ and $z$ satisfy the relations in (2.14).
@@
Now, we need to compute the series $S_{2p-1,1}^{{}^{\prime}}\left(y\right)$. According to Lemma 2.5, there exist polynomials ${\rm q}_{2n-1}(x)\in\mathbb{Z}[x]$ such that the Maclaurin series of ${\rm sn}(u)$ has the form
@@
From [8], we recall the Fourier series expansions of Jacobi elliptic functions ${\rm sn}(u)$
@@
Setting $q\equiv q\left(x\right)=e^{-y}$, we then have
${\rm sn}(u)$ $=\displaystyle\frac{2\pi}{Kk}\displaystyle\sum_{n=0}^{\infty}{\displaystyle\frac{q^{n+\tfrac{1}{2}}\left(2n+1\right)^{2j+1}}{1-q^{2n+1}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2K}\right)^{2j+1}}{\left(2j+1\right)!}}$
$=\displaystyle\frac{\pi}{Kk}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{2K}\right)^{2j+1}}{\left(2j+1\right)!}}S_{2j+1,1}^{\prime}\left(y\right).$ (4.48)
Comparing the coefficients of $u^{2m+1}$ in (4.46) and (4), we can deduce
@@
Differentiating (4.49) with respect to $y$ yields
@@
Using the relation $\frac{dx}{dy}=-x(1-x)z^{2}$ (see [5, P. 120, Entry. 9(i)]), we obtain
@@
Finally, combining (4.45), (4.49), (4.51) and (4) with elementary calculations yields the desired result. ∎
Using Mathematica, we can easily compute the following using the formulas in Theorem 4.1.
@@
where $E$ is the complete elliptic integrals of the second kinds. Applying (4.46), we deduce the power series expansion of the function $\mathrm{sn}^{2}(u)$
@@
Differentiating (4.55) with respect to $y$ gives
@@
From [25, Cor.1], we also obtain
@@ 5 Evaluations of Mixed Berndt-type Integrals
We assume that $x_{0},y_{0}$ and $z_{0}$ are all known. For $\Omega\left(x_{0},\exp\left(-y_{0}\right),z_{0},z_{0}^{\prime}\right)=0$, considering the transformation (2.18)
@@
According to Lemma 2.1, we know that $\text{q}_{2p-1}(x)=\displaystyle\sum^{p-1}_{j=0}a_{j}x^{j}$. Applying Theorem 4.1 and transformation (5.60), one obtains
@@
Setting $x_{0}=1/2$ in (5.63) yields $y_{0}=\pi$ via (2.14). All coefficients preceding $z_{0}^{2p+2}$ and $z_{0}^{2p+1}z_{0}^{\prime}$ are rational numbers. When $x_{0}=1/2$, the coefficient of $z_{0}^{2p+1}z_{0}^{\prime}$ is
@@
In fact, Lemma 2.5 indicates that $a_{j}=a_{p-1-j}$ for $j=0,\ldots,p-1$. From (2.24), we can directly conclude that
@@
$=a_{j}\left\{\left(\sqrt{2}-1\right)^{4j-2p+2}+\left(\sqrt{2}+1\right)^{2p-2}\left(\sqrt{2}-1\right)^{4\left(p-1-j\right)}\right\}$
$=a_{j}\left\{\left(\sqrt{2}-1\right)^{4j-2p+2}+\left(\sqrt{2}+1\right)^{4j-2p+2}\right\}\in\mathbb{Z},\quad(j=1,\ldots,[(p-1)/2]).$ (5.65)
where $\cdot$] is the floor function. Further, for $j=0,\ldots,p-1$, we obtain
@@
$\quad\quad=\displaystyle\frac{\left(-1\right)^{p}}{2^{3p+1}}\sqrt{2}\left(1+\sqrt{2}\right)^{2p-4}\displaystyle\sum_{j=0}^{p-1}{a_{j}}\left\{\left(2-\sqrt{2}\right)p-\left(4j+2\right)\right\}\left(\sqrt{2}-1\right)^{4j-2}.$ (5.73)
According to (2.25), (2.24) and $a_{j}=a_{p-1-j},j=0,\ldots,p-1$, we have
$\sqrt{2}\left(1+\sqrt{2}\right)^{2p-4}\left\{\begin{array}[]{c}a_{j}\left\{\left(2-\sqrt{2}\right)p-2\left(2j+1\right)\right\}\left(\sqrt{2}-1\right)^{4j-2}+\\ a_{p-1-j}\left\{\left(2-\sqrt{2}\right)p-2\left(2p-2j-1\right)\right\}\left(\sqrt{2}-1\right)^{4p-4j-6}\\ \end{array}\right\}$ (5.76)
$=a_{j}\left\{\begin{array}[]{c}\left\{2\left(\sqrt{2}-1\right)p-2\sqrt{2}\left(2j+1\right)\right\}\left(\sqrt{2}+1\right)^{2p-4j-2}\\ +\left\{2\sqrt{2}\left(2j+1\right)-2\left(\sqrt{2}+1\right)p\right\}\left(\sqrt{2}-1\right)^{2p-4j-2}\\ \end{array}\right\}$ (5.79)
$=2a_{j}\left\{\begin{array}[]{c}\sqrt{2}\left(2j+1\right)\left\{\left(\sqrt{2}-1\right)^{2p-4j-2}-\left(\sqrt{2}+1\right)^{2p-4j-2}\right\}\\ +p\left\{\left(\sqrt{2}+1\right)^{2p-4j-3}-\left(\sqrt{2}-1\right)^{2p-4j-3}\right\}\\ \end{array}\right\}\in\mathbb{Z},$ (5.82)
where $j=1,\ldots,[(p-1)/2]$. Combining (5.66) and (5.82) gives
@@
We note the $n$-th derivative of $z$, as given in (2.14), is
$\displaystyle\frac{\mathrm{d}^{n}z}{\mathrm{d}x^{n}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}}{n!}\,_{2}F_{1}\left(\displaystyle\frac{1}{2}+n,\displaystyle\frac{1}{2}+n,1+n;x\right).$ (5.84)
To evaluate this derivative at $x_{0}=\displaystyle\frac{1}{2}$, we apply the known identity ([2])
$\,_{2}F_{1}\left(a,b,\displaystyle\frac{a+b+1}{2};\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma\left(\displaystyle\frac{1}{2}\right)\Gamma\left(\displaystyle\frac{a+b+1}{2}\right)}{\Gamma\left(\displaystyle\frac{a+1}{2}\right)\Gamma\left(\displaystyle\frac{b+1}{2}\right)}.$ (5.85)
By substituting $a=b=\displaystyle\frac{1}{2}+n$ into (5.85) and setting $x_{0}=\displaystyle\frac{1}{2}$ in (5.84), we obtain
$\displaystyle\frac{d^{n}z}{dx^{n}}\mid_{x=x_{0}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}\sqrt{\pi}}{\Gamma^{2}\left(\displaystyle\frac{n}{2}+\displaystyle\frac{3}{4}\right)}.$ (5.86)
Thus, we evaluate $y_{0}=\pi,z_{0}=\displaystyle\frac{\Gamma^{2}}{2\pi^{3/2}}$, $z_{0}^{\prime}=\displaystyle\frac{4\pi^{1/2}}{\Gamma^{2}}$ by (2.14) and (5.86) , then substituting them into (5.83) yields
@@
In Theorem 3.1, letting a=4m2a=4m-2 and $4m$, we have
@@
Remark 5.2.
In particular, by using the methods of the proof in [6, Thm 4.1], we consider
$\mathscr{Q}=\displaystyle\lim_{R\rightarrow\infty}\displaystyle\int_{D_{R}}{\displaystyle\frac{\mathrm{d}z}{\left(\sinh z+i\sin z\right)\left(\cosh z+\cos z\right)}},$
where $D_{R}$ is the same contour that we considered above, except that now the contour is indented at the origin by a quarter-circle $\Gamma_{\epsilon}$ of radius $\epsilon$.
On $\Gamma_{\epsilon}$, set $z=\epsilon e^{i\theta}$, $\frac{1}{2}\pi\geq\theta\geq 0$, and the contour is shown in the following figure:
It can be obtained by the residue method that
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\pi}{8}-\displaystyle\frac{\pi}{2}\bar{C}_{0,2}\left(\displaystyle\frac{\pi}{2}\right),$ (5.90)
and by using Weierstrass elliptic functions [13, 14] to obtain closed forms of $\bar{C}_{0,2}\left(\displaystyle\frac{\pi}{2}\right)$
$\bar{C}_{0,2}\left(\displaystyle\frac{\pi}{2}\right)=\displaystyle\frac{1}{2}-\displaystyle\frac{\Gamma^{4}}{4\pi^{5}}.$ (5.91)
Substituting (5.91) into (5.90) , we can find that
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\Gamma^{4}}{32\pi^{2}}-\displaystyle\frac{\pi}{8}.$ (5.92)
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{8}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=\displaystyle\frac{33\Gamma^{20}}{2097152\pi^{6}}-\displaystyle\frac{63\Gamma^{16}}{262144\pi^{4}}.\$
@@
$\quad\quad\times\left[\begin{array}[]{c}z_{0}\left(1+\sqrt{x_{0}}\right)\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}\mid_{4\sqrt{x_{0}}/\left(1+\sqrt{x_{0}}\right)^{2}}\\ +2p\displaystyle\frac{1+x_{0}-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\mathrm{Q}_{2p-2}\left(\displaystyle\frac{1+x_{0}-2\sqrt{x_{0}}}{1+x_{0}+2\sqrt{x_{0}}}\right)\left\{\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}\right\}\\ \end{array}\right]$ (5.101)
@@
$\quad\quad\times\left\{\sqrt{x_{0}}\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{5}}{1-x_{0}}\displaystyle\frac{z_{0}^{\prime}}{2}+\displaystyle\frac{\left(1+\sqrt{x_{0}}\right)^{4}}{1-x_{0}}\displaystyle\frac{z_{0}}{4}\right\}+\displaystyle\frac{\left(-1\right)^{p-1}\left(1-x_{0}\right)^{2}{z_{0}}^{2p+2}}{2^{2p-2}}\sqrt{x_{0}}\left(1+\sqrt{x_{0}}\right)^{2p-4}$
$\quad\quad\times\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\left\{\left(1-x\right)\mathrm{Q}_{2p-2}\left(1-x\right)\right\}\mid_{4\sqrt{x_{0}}/\left(1+\sqrt{x_{0}}\right)^{2}}.$ (5.102)
@@
$\mathrm{Q}_{2p-2}\left(x\right)=\displaystyle\sum_{j=0}^{p-2}{b_{j}x^{j}}\quad(b_{j}\in\mathbb{Z}),$ (5.104)
@@
$=b_{j}\left\{\left(\sqrt{2}-1\right)^{4j-2p+4}+\left(\sqrt{2}+1\right)^{4j-2p+4}\right\}\in\mathbb{Z}\quad(j=0,\ldots,p-2).$ (5.105)
@@
$=\displaystyle\frac{p}{4}\left[\left(\sqrt{2}+1\right)^{2p-4j-5}-\left(\sqrt{2}-1\right)^{2p-4j-5}\right]$
$\quad+\displaystyle\frac{\sqrt{2}}{2}\left(j+1\right)\left[\left(\sqrt{2}-1\right)^{2p-4j-4}-\left(\sqrt{2}+1\right)^{2p-4j-4}\right]\in\mathbb{Q},$ (5.114)
where $j=0,\ldots,[(p-2)/2]$. Combining (5.106) with (5.113), we derive that
@@
In Theorem 3.3, letting $a=4m$ and 4m24m-2, we have